### v1

# SpatialCrafter: Unleashing the Imagination of Video Diffusion Models for Scene Reconstruction from Limited Observations

Abstract
Novel view synthesis (NVS) boosts immersive experiences in computer vision and graphics. Existing techniques, though progressed, rely on dense multi-view observations, restricting their application. This work takes on the challenge of reconstructing photorealistic 3D scenes from sparse or single-view inputs. We introduce SpatialCrafter, a framework that leverages the rich knowledge in video diffusion models to generate plausible additional observations, thereby alleviating reconstruction ambiguity. Through a trainable camera encoder and an epipolar attention mechanism for explicit geometric constraints, we achieve precise camera control and 3D consistency, further reinforced by a unified scale estimation strategy to handle scale discrepancies across datasets. Furthermore, by integrating monocular depth priors with semantic features in the video latent space, our framework directly regresses 3D Gaussian primitives and efficiently processes long-sequence features using a hybrid network structure. Extensive experiments show our method enhances sparse view reconstruction and restores the realistic appearance of 3D scenes.
Figure 1: Overview. Our method, SpatialCrafter, generates additional novel views of a scene from few inputs by leveraging camera trajectory-guided video diffusion models, which alleviates the ambiguity of sparse view reconstruction and robustly reconstructs the scene from the generated video frames. It demonstrates impressive performance in both indoor and outdoor scenes, while also exhibiting generalization capabilities in real scenes and synthetic images. More visualization results can be found in the supplementary material.
1 Introduction
Novel View Synthesis (NVS), as a key technology in computer vision and graphics, provides crucial support for immersive experiences in fields such as video games and mixed reality. Although neural reconstruction techniques have made significant progress in recent years [25, 58, 62, 41, 38], these methods typically rely on dense multi-view observation data, which face numerous limitations in practical applications. Therefore, this paper focuses on a more practically valuable challenge: how to achieve high-quality 3D scene reconstruction and synthesize realistic novel views from sparse or even single-view observation.
The complexity of this problem stems from three main aspects: first, the diversity and complexity of real-world scene make precise modeling extremely challenging; second, occluded regions in sparse viewpoints are difficult to recover; and finally, limited geometric cues constrain the accuracy of reconstruction. Compared to traditional methods that depend on numerous viewpoints, achieving high-quality reconstruction from sparse views requires stronger prior knowledge and more efficient reasoning capabilities, which is the core challenge of this problem. To address these challenges, some methods [29, 81, 26, 39, 18, 77] constrain the optimization process by introducing regularization terms, but these methods need to be optimized for each scene, are not generalizable, and are difficult to cope with complex scenarios. Recently, some feed-forward methods [8, 9, 56, 48, 22, 53, 73, 74] have achieved generalizable sparse view scene reconstruction by directly predicting 3D Gaussian primitives for each pixel. However, such methods produce severe artifacts and implausible images when reconstructing occluded regions and applying them to the extrapolation setting. To address this challenge, some methods [34, 60, 43, 61] have introduced the priors from the diffusion model to achieve sparse or single-view NVS, but they perform poorly on scene-level data and lack precise pose control and 3D consistency. This is due to the fact that low-dimensional camera information such as Euler angles and extrinsic struggle to provide comprehensive control signals to the generative models.
In this paper, we propose a framework named SpatialCrafter for scene reconstruction and novel view synthesis based on sparse or single-view inputs. Our innovation lies in leveraging the rich physical world knowledge embedded in video diffusion models to provide plausible additional observations for scene reconstruction, thereby effectively reducing the problem complexity. Specifically, we first focus on enhancing precise camera control and 3D consistency in generated videos. To achieve this, we parameterize camera settings using ray embeddings or metric depth-warped images and incorporate them into the video diffusion model via a trainable camera encoder. Furthermore, we propose an epipolar attention mechanism that improves 3D consistency between video frames through explicit geometric constraints. Previous methods [60, 20] often experience performance degradation and difficulty generating large-motion videos when trained on multiple scene datasets. We identified that these issues primarily stem from scale ambiguity in scene datasets. To overcome this, we introduce a unified scale estimator to calibrate scene dataset, enabling effective joint training across multiple datasets.
Although we can generate visually coherent video sequences, relying solely on generated frames to reconstruct general scenes often leads to suboptimal solutions, particularly in large-scale outdoor environments and stylized scenes. To address this, we propose incorporating monocular depth priors with rich semantic features extracted from the video latent space. Leveraging these latent features, our method directly regresses 3D Gaussian primitives of the scene via a feed-forward network. Furthermore, to efficiently handle long-sequence feature interactions, we design a hybrid architecture that integrates Mamba blocks with Transformer blocks. Experiments show that our method not only improves the sparse-view reconstruction, but also accurately restores the appearance of 3D scenes, especially when extrapolating from a single view and when there are few sparse view overlaps. In conclusion, our key contributions can be summarized as follows:
  • We introduce a framework that effectively utilizes the physical-world knowledge embedded in video diffusion models to provide additional plausible observations for sparse-view scene reconstruction, thus reducing the ambiguity of sparse view scene reconstruction.
  • To address the scale ambiguity problem that occurs in joint training across datasets, we develop a unified scale estimation approach for trajectory calibration. This solves the performance degradation problem, thus enabling effective multi-dataset training.
  • We combine monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives through a feed-forward manner. Meanwhile, we propose a hybrid architecture integrating Mamba blocks with Transformer blocks to efficiently handle long-sequence feature interactions.
2 Related Work
Figure 2: Overview of our pipeline. Our generative reconstruction pipeline consists of two parts: camera-controlled video generation, and then reconstruction. First, we set the exploration path based on the input views. In the video generation module, ray embedding is used to parameterize the camera information, and a cross-attention mechanism with epipolar constraints is introduced to improve the 3D consistency of the generated video. The robust 3D scene reconstruction pipeline integrates monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives through a feed-forward manner.
2.1 Sparse-View Scene Reconstruction
NeRF [58] and Gaussian Splatting  [25] have achieved photorealistic representations of 3D scenes, but require optimization on densely collected image sets for each scene. To address this problem, some approaches [29, 18, 81, 67] focus on constraining the optimization process of 3D representations via manually designed regularization. Some other methods [23, 8, 9, 52, 65, 68] are trained on large-scale datasets and directly predict the 3D representation in a feed-forward manner. Recently, some approaches [14, 33, 48, 30] based on end-to-end 3D reconstruction models [56, 28], have achieved efficient and high-quality 3D reconstruction. However, these methods cannot handle occluded areas, and when reconstruction fails (e.g., in the case of large viewpoint changes), the results of subsequent novel view synthesis can be significantly compromised.
2.2 Diffusion-based Novel View Synthesis
Since text-to-image diffusion models [42] contain rich natural image priors, some methods [40, 59] directly optimize 3D representations [58, 25] via score distillation sampling [40]. Several approaches [34, 43, 54, 7] enable zero-shot novel view synthesis by fine-tuning the image and pose-conditioned generative models on large-scale 3D object [11] or scene [80, 54] datasets. However, they still have difficulty synthesizing consistent novel views. To solve this, some approaches [35, 78, 16, 37, 45, 61, 6] generate multi-views simultaneously and model the correlation between multiple views with ground-truth poses. Some recent methods [36, 30] proposed to use large-scale unposed images to train NVS methods. In addition, some methods [10, 76, 15, 70, 46, 13] utilize depth-based warping to synthesize novel views and employ the T2I model to inpaint the warped images. However, the novel views generated by these methods tend to suffer from artifacts and cumulative errors, limiting their use in generalized scenes.
2.3 Controllable Video Generation
While recent text-to-video diffusion models have achieved remarkable progress, they inherit the controllability limitations of their text-to-image counterparts, often requiring additional conditioning mechanisms to align generated videos with user intent. Several works [57, 19, 24, 66] have been carried out to introduce a variety of conditions into video generation models. Recently, attention has focused on controlling the camera motion of the video. Some approaches [19, 51] employs motion-specific training for predefined camera movements, though this framework struggles with complex trajectory synthesis. Some methods [60, 20, 63] enable more complex camera control by parameterizing the camera trajectory and injecting it into a pre-trained video diffusion model via a trainable camera encoder. MotionCtrl [60] controls camera and object motions independently through learnable adapters, while CameraCtrl [20] introduces Plücker embeddings for accurate 6-DoF camera pose representation. AC3D [1] optimizes the pose adjustment scheme to accelerate convergence and limits the injection of camera features to specific locations. ViewCrafter [71] and ReconX [33] pushes this direction by enabling pixel-accurate view compositing through point-based rendering formulation.
3 Method
This section begins by explaining our camera-controlled video generation model (Section 3.1), which employs ray embeddings for precise pose control. Then, we describe how to use epipolar geometry constraints to enhance 3D consistency in the video frames. Next, we introduce ourscene reconstruction pipeline (Section 3.2), which aims to perform stable 3D reconstruction based on the generated video latents. These innovations collectively improve the accuracy and consistency of 3D reconstructions from generated video.
3.1 Camera-Conditioned Video Generation
Scale Alignment. While recent methods [60, 20] have achieved promising results in novel view synthesis, their reliance on ground-truth camera extrinsic from Re10K [79] during inference inherently limits their applicability to in-the-wild scenarios. Moreover, the metric-scaled poses in Re10K create alignment challenges when jointly training with other datasets (e.g., DL3DV [31]) where poses are estimated using COLMAP, primarily due to scale ambiguity issues. To address this issue, we introduced a metric depth estimation model [5] to achieve scale alignment of camera poses between different datasets. During the data preprocessing stage, we align the metric 3D point cloud predicted by the model with the sparse point cloud generated using COLMAP to calculate the scale factor, thereby obtaining absolute-scale camera extrinsics.
Camera Injection. Inspired by recent methods [20, 72, 47], we chose to use ray embeddings [20] or the depth warping frames [71] to represent the camera information. This over-parameterized approach [72] provides more accurate pose information to the video generator than other methods. Specifically, for a ray defined by the origin $o\in\mathbb{R}^{3}$ and the normalized direction $d\in\mathbb{R}^{3}$, we can represent it as $(o\times d,d)\in\mathbb{R}^{6}$. Given the camera parameters, the direction of the ray $d^{\prime}$ corresponding to the pixel coordinates $(u,v)$ can be calculated as $d^{\prime}=\mathbf{R}\mathbf{K}^{-1}(u,v,1)^{\top}+\mathbf{T}$. For depth warping frames, we leverage the pretrained depth estimation model [55] to map the reference image into a 3D point cloud. Subsequently, we render novel views at specified camera poses by projecting this point cloud via the target camera parameters. Afterwards, we use a trainable conditional encoder to inject the camera information into the video model instead of fine-tuning all model parameters. The training objective is:
$\mathcal{L}=\mathbb{E}_{z,z_{0},\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},t,\phi({c}))\right\|_{2}^{2}\right]$ (1)
where $\boldsymbol{P}$ represents the ray embeddings, $z_{0}$ denotes the latent embeddings of the reference frame, and $t$ indicates the timestamps.
Epipolar Feature Aggregation. Camera embedding enhances control but 3D video consistency remains challenging due to dense self-attention allowing unrestricted cross-frame pixel interactions. We address this by incorporating epipolar geometric constraints, and aggregating features along epipolar lines. For a pixel $\boldsymbol{p}=(u,v)$ in frame $i$, its corresponding epipolar line $\boldsymbol{l}_{ik}$ in frame $k$ is computed as $\boldsymbol{l}_{ik}(\boldsymbol{p})=\boldsymbol{F}_{ik}\cdot\tilde{\boldsymbol{p}}$, where $\tilde{\boldsymbol{p}}=(u,v,1)^{\top}$ are the homogeneous coordinates, and $\boldsymbol{F}_{ik}\in\mathbb{R}^{3\times 3}$ is the fundamental matrix. The fundamental matrix is decomposed as $\boldsymbol{F}_{ik}=\boldsymbol{K}_{k}^{-\top}\boldsymbol{E}_{ik}\boldsymbol{K}_{i}^{-1}$, where $\boldsymbol{K}_{i},\boldsymbol{K}_{k}\in\mathbb{R}^{3\times 3}$ are the camera intrinsics, and $\boldsymbol{E}_{ik}$ is the essential matrix. We then convert epipolar lines into attention masks by computing per-pixel distances and applying a threshold, restricting attention to geometrically valid regions.
Sparse-View Setting. To adapt to the sparse view input setting, we formulate the task as a video frame interpolation problem conditioned on the given start and end frames. To maximally preserve the priors from the pre-trained models [4], we inject boundary frame conditions in both latent and semantic space. Specifically, we combine latent features of the first and last frames with their noisy latent representations, and then concatenate the extracted CLIP embeddings from both frames for cross-attentional feature injection. The training objective is:
$\mathcal{L}=\mathbb{E}_{z,z_{0},z_{n}\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},z_{n},t,\boldsymbol{P})\right\|_{2}^{2}\right]$ (2)
where $z_{n}$ and $z_{0}$ are the latent embeddings of the final and initial frames, respectively. During the inference phase, we use DUSt3R to estimate the extrinsic and intrinsic parameters for the input views.
Figure 3: Qualitative comparison of generated videos. Compared to other benchmark methods, the videos generated by our method have better foreground object-background consistency, while also being able to generate videos with larger motion amplitude.
3.2 Video-based Scene Reconstruction
Motivation. There are several challenges in reconstructing scenes directly from generated videos. First, the limited number of generated frames makes it difficult to capture complete scene information. In addition, the diverse styles of video frames can pose a challenge to traditional reconstruction techniques. Furthermore, the generated video frames may contain imperfect or low-quality areas, making the reconstruction process unstable. Moreover, In stylized scenes, previous methods often yield poor results, as recovering poses and sparse point clouds from video frames is difficult, leading to many artifacts. To address this, a powerful reconstruction module specifically designed to handle the generated videos is required.
[…]
RealEstate10K
MotionCtrl [60] 22.65 230.12 0.234 0.798 0.299 14.72 0.404
CameraCtrl [20] 21.48 188.21 0.054 0.127 0.230 17.33 0.516
ViewCrafter [71] 20.96 204.18 0.055 0.153 0.215 18.95 0.503
Ours 18.25 183.25 0.052 0.103 0.207 19.21 0.523
Tanks-and-Temples
MotionCtrl [60] 30.25 290.38 0.838 1.505 0.315 14.64 0.388
CameraCtrl [20] 24.41 244.82 0.118 0.294 0.285 15.38 0.469
ViewCrafter [71] 22.50 231.35 0.126 0.307 0.247 16.24 0.508
Ours 20.17 192.52 0.097 0.175 0.228 16.12 0.501
DL3DV
MotionCtrl [60] 25.65 249.51 0.473 1.118 0.312 14.38 0.386
CameraCtrl [20] 22.76 233.54 0.095 0.238 0.262 16.33 0.489
ViewCrafter [71] 20.59 211.24 0.093 0.244 0.242 17.12 0.521
Ours 18.21 171.52 0.063 0.134 0.225 17.02 0.537
Table 1: Quantitative comparison to the camera conditioned video generation method on RealEstate10K, DL3DV, and Tanks-and-Temples dataset.
Latent Feature Fusion. Given input video latents $\mathbf{z}\in\mathbb{R}^{T\times H\times W\times C}$ and camera poses encoded as ray embeddings $\mathbf{p}\in\mathbb{R}^{T\times H\times W\times 6}$, we transform them into token sequences through a patchification process. Specifically, we apply spatial patchification to latent features to obtain latent tokens $\mathbf{z}_{t}\in\mathbb{R}^{N\times d_{z}}$, while ray embeddings undergo 3D-patchification to produce pose tokens $\mathbf{p}_{t}\in\mathbb{R}^{N\times d_{p}}$. To incorporate explicit geometric guidance, we leverage monocular depth estimation within our pipeline. For each input RGB video frame, we estimate dense depth maps $\mathbf{D}\in\mathbb{R}^{T\times H\times W\times 1}$, which are processed by a dedicated depth encoder to yield depth tokens $\mathbf{d}_{t}\in\mathbb{R}^{N\times d_{d}}$. The three token sets are channel-wise concatenated to form a unified representation $\mathbf{x}=[\mathbf{z}_{t};\mathbf{p}_{t};\mathbf{d}_{t}]\in\mathbb{R}^{N\times(d_{z}+d_{p}+d_{d})}$, which is then linearly projected to a lower-dimensional space $\mathbf{x}^{\prime}\in\mathbb{R}^{N\times d}$ before being fed into a sequence of mamba and transformer blocks. Mamba [17] achieves the same token sequence processing functionality as Transformer, but reduces computational complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L)$, making it particularly suitable for dense reconstruction tasks. In our Mamba blocks, we implement bi-directional scanning across the token sequence. First, we compute state parameters $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times d}$ from the input tokens using a linear projection. Then, we execute the State Space Model (SSM) in both forward ($\mathbf{y}_{f}$) and backward ($\mathbf{y}_{b}$) directions. The final output of each Mamba block is computed as $\mathbf{y}=\mathbf{y}_{f}+\mathbf{y}_{b}$ followed by a final linear transformation.
Gaussian Decoding. We design a lightweight decoder module that efficiently transforms output feature tokens into per-pixel Gaussian parameters. The decoding module consists of 3D-DeConv layers, which generates refined Gaussian feature map $G\in\mathbb{R}^{(T\times H\times W)\times 12}$. This 12-channel representation precisely encodes the complete set of Gaussian parameters: RGB color (3 channels), scale factors (3 channels), rotation quaternion (4 channels), opacity (1 channel), and ray distance (1 channel). The final output of the model is the merge of 3D Gaussians from all input video frames.
Training Objective. During training, we render images from predicted Gaussians using randomly selected supervision views. Our approach employs a composite loss that integrates three components:
$\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mse}+\lambda_{2}\mathcal{L}_{perc}+\lambda_{3}\mathcal{L}_{depth},$ (3)
where $\mathcal{L}_{mse}$ represents pixel-wise mean squared error, $\mathcal{L}_{perc}$ denotes perceptual loss, and $\mathcal{L}_{depth}$ enforces depth consistency. This formulation jointly optimizes for photometric accuracy, high-level perceptual fidelity, and geometric coherence.
4 Experiment
4.1 Experiment settings
Datasets. To comprehensively capture the underlying distribution of real-world scenarios, we trained our video diffusion model on three diverse datasets: RealEstate-10K [79], ACID [32], and DL3DV-10K [31]. The Re10K dataset from YouTube comprises 67,477 training and 7,289 testing indoor and outdoor camera trajectories. The ACID dataset focuses on natural landscapes, with 11,075 training and 1,972 testing scenes. The DL3DV-10K dataset is a large-scale collection featuring 10,510 scenes captured under controlled, standardized conditions.
Figure 4: Qualitative comparison on Re10K [79]. Compared to baselines, we obtain superior reconstruction from limited overlap, and enhanced geometry reconstruction in non-overlapping regions.
Figure 5: Qualitative comparison on DL3DV and Tank-and-Temples dataset. Our method reconstructs better than baselines, even with limited image overlap and in non-overlapping regions.
Implementation Details. Our video generation model is based on SVD [3], an image-video diffusion model based on UNet. In our experiments, we use a relative camera system in which all camera poses are converted to poses relative to the first frame. The camera in the first frame is located at the world origin. In the first stage, we train the model at a resolution of $320\times 512$ for 50,000 iterations with the frame length set to 25. Subsequently, we train on $576\times 1024$ for 10,000 iterations to adapt to high resolution. The learning rate is set to $51e-$ 5 with a warmup of 1,000 steps, using the Adam optimizer. We chose Lightning as the training framework, using mixed-precision fp16 and DeepSpeed ZeRO-2. We trained the proposed method and its variants on 16 NVIDIA A800 GPUs with a batch size of 32. During inference, we adopt DDIM sampler [49] with classifier-free guidance [21].
Progressive Training. Controlling a video generation model to generate arbitrary motion trajectories remains a challenging task. To enable robust arbitrary trajectory generation, we employ a three-stage curriculum learning strategy. The model first trains on smooth camera motions with small temporal intervals, then progressively adapts to more complex motion patterns through linear interval scheduling, and finally incorporates random sampling intervals. This gradual transition from simple to complex trajectories proves crucial for high-quality video generation with arbitrary trajectories.
3-view 6-view 9-view
Method PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓
Mip-NeRF360
Zip-NeRF [2] 12.77 0.271 0.705 13.61 0.284 0.663 14.30 0.312 0.633
ReconFusion [61] 15.50 0.358 0.585 16.93 0.401 0.544 18.19 0.432 0.511
CAT3D [16] 16.62 0.377 0.515 17.72 0.425 0.482 18.67 0.460 0.460
ReconX [33] 17.16 0.435 0.407 19.20 0.473 0.378 20.13 0.482 0.356
ViewCrafter [71] 14.51 0.315 0.682 15.87 0.336 0.665 16.45 0.348 0.655
Ours 17.32 0.439 0.418 19.42 0.510 0.371 20.19 0.536 0.345
DTU
Zip-NeRF [2] 9.18 0.601 0.383 8.84 0.589 0.370 9.23 0.592 0.364
FSGS [81] 17.34 0.818 0.169 21.55 0.880 0.127 24.33 0.911 0.106
ReconFusion [61] 20.74 0.875 0.124 23.62 0.904 0.105 24.62 0.921 0.094
CAT3D [16] 22.02 0.844 0.121 24.28 0.899 0.095 25.92 0.928 0.073
ViewCrafter [71] 15.63 0.522 0.383 15.03 0.525 0.408 14.92 0.487 0.452
Ours 27.92 0.879 0.103 28.82 0.895 0.082 29.03 0.905 0.065
LLFF
Zip-NeRF [2] 17.23 0.574 0.373 20.71 0.764 0.221 23.63 0.830 0.166
FSGS [81] 20.31 0.652 0.288 24.20 0.811 0.173 25.32 0.856 0.136
ReconFusion [61] 21.34 0.724 0.203 24.25 0.815 0.152 25.21 0.848 0.134
CAT3D [16] 21.58 0.731 0.181 24.71 0.833 0.121 25.63 0.860 0.107
ViewCrafter [71] 17.73 0.521 0.332 17.43 0.512 0.345 17.33 0.488 0.351
Ours 22.04 0.741 0.165 25.11 0.814 0.124 25.95 0.838 0.103
Table 2: Quantitative comparison of sparse view 3D reconstruction on Out-of-Domain Datasets.
RealEstate10K ACID
Method PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓
pixelNeRF [69] 20.43 0.589 0.550 20.97 0.547 0.533
GPNR [50] 24.11 0.793 0.255 25.28 0.764 0.332
AttnRend [12] 24.78 0.820 0.213 26.88 0.799 0.218
MuRF [64] 26.10 0.858 0.143 28.09 0.841 0.155
pixelSplat [8] 25.89 0.858 0.142 28.14 0.839 0.150
MVSplat [9] 26.39 0.869 0.128 28.25 0.843 0.144
GS-LRM [73] 28.10 0.892 0.114 - - -
DepthSplat [65] 24.23 0.790 0.217 - - -
ReconX [33] 28.31 0.912 0.088 28.84 0.891 0.101
ViewCrafter [71] 24.22 0.788 0.218 23.48 0.660 0.299
Ours 28.35 0.862 0.121 28.93 0.899 0.116
Table 3: Quantitative comparison of two-view sparse view reconstruction methods on RealEstate10K [79] and ACID [32] dataset.
4.2 Comparisons
4.2.1 Controllable Video Generation
Benchmark and Metrics. We evaluate our video generation model against three baselines [60, 20, 71] on three benchmark datasets: RealEstate10K [79] (500 randomly selected videos with first frame as image condition and subsequent frames at stride 3 for pose guidance), DL3DV-140 [31] (500 video clips at stride 2), and Tanks-and-Temples [27] (100 sequences at stride 4 from 14 scenes with COLMAP-annotated poses) for out-of-domain generalization testing.
Metrics. Our evaluation framework employs multiple metrics to assess performance: visual quality measured through Fréchet Video Distance (FVD) and Fréchet Inception Distance (FID); camera control precision quantified by rotation error ($R_{err}$) and translation error ($T_{err}$) computed from camera poses extracted via COLMAP and normalized relative to the first frame; and visual consistency evaluated using PSNR, SSIM, and LPIPS [75] between generated and ground-truth views. To ensure fair comparison across methods with varying output capabilities, we restrict visual similarity assessment to the first 14 frames, as generated content typically diverges progressively from the conditional single-view input as the scene extends.
Comparison. Table 1 shows that our method excels in qualitative results, outperforming other methods in terms of visual quality, pose control, and 3D consistency. Figure 3 further demonstrates our advantage in 3D consistency and visual quality. This performance improvement is attributed to the introduction of a camera parameterization method and an epipolar attention module, which respectively enhance pose control ability and the 3D consistency of generated videos. In addition, training on a diverse scene dataset significantly improves the generalization ability of the model.
Small Large
[…]
pixelNeRF [69] 18.417 0.601 0.526 20.869 0.639 0.458
AttnNeRF [12] 19.151 0.663 0.368 25.897 0.845 0.229
pixelSplat [8] 20.263 0.717 0.266 27.151 0.879 0.122
MVSplat [9] 20.353 0.724 0.254 27.408 0.884 0.116
DUSt3R [56] 14.101 0.432 0.468 16.427 0.453 0.402
CoPoNeRF [22] 17.393 0.585 0.462 20.464 0.652 0.358
Ours 22.514 0.784 0.213 27.411 0.913 0.109
Table 4: Quantitative comparison on the RealEstate10K dataset. Small and Large refer to input images with low and high overlap ratios, respectively. Greater overlap means greater temporal coherence between adjacent images.
4.2.2 Sparse View Reconstruction
Benchmark and Metrics. To comprehensively evaluate the performance of our method, we compare it with two categories of reconstruction methods: (1) two-view feed-forward reconstruction methods [8, 12, 9, 73, 65, 50], and (2) optimization-based sparse-view reconstruction methods [14, 16, 71, 33]. The evaluation employs two types of test sets: in-domain test sets (from RealEstate10K [79] and Acid [32]) and out-of-domain test sets. To ensure a fair comparison, the out-of-domain test set is the same as that used in previous work [61, 16]. For evaluating scene reconstruction quality, we adopt three metrics: PSNR, SSIM, and LPIPS [75]. Additionally, we categorize the evaluation into two groups based on the overlap between input views, with detailed overlap estimation methods provided in the supplementary materials. The evaluation protocol follows a consistent procedure: first, we reconstruct the 3D scene based on the input images, then render from novel views, and compute the metrics between the rendered images and the reference images.
Comparison. As shown in Tables 2-4 and Figures 5-5, our method performs well on various evaluation metrics, particularly in challenging scenarios with occlusions and limited view overlap. This improvement can be attributed to our approach that utilizes scene priors and sparse input views to generate helpful additional observations, thereby enhancing visual correlations between different viewpoints and helping to reduce the complexity of sparse-view reconstruction. Compared to the method [71] conditioning on point-based rendering, our approach shows advantages in reconstructing scenes with thin structures, leading to improved 3D reconstruction results.
Figure 6: Qualitative comparisons of novel views rendered from scenes reconstructed using other 3D generation methods.
Tanks-and-Temples DL3DV
[…]
LucidDreamer [10] 14.552 0.364 0.415 15.126 0.452 0.431
ZeroNVS [43] 14.734 0.381 0.480 15.163 0.464 0.466
MotionCtrl [60] 15.322 0.426 0.404 16.889 0.528 0.392
ViewCrafter [71] 21.286 0.654 0.193 21.373 0.686 0.244
Ours 22.331 0.742 0.213 22.782 0.791 0.202
Table 5: Quantitative comparison of single-view novel view synthesis on DL3DV [31] and Tank-and-Temples [27] benchmarks.
4.2.3 Single View 3D Generation
Benchmark and Metrics. In single-view generation, we compare our method with several generative methods [43, 60, 71, 44] on the Tank-and-Temples [27] and DL3DV-140 [31] datasets. Noted that we begin by reconstructing the 3D scene from the given images, followed by calculating the metrics using renderings from novel views. Similarly, we used the PSNR, SSIM, and LPIPS [75] metrics to evaluate our results. Evaluation in this underconstrained setting is challenging, as multiple 3D scenes can produce consistent generations for a given view. Thus, we measure metrics using only temporally adjacent frames to the conditional image, 14 sampled frames and poses captured after the conditional image.
Comparison. As shown in Table 5, our method consistently achieves superior performance on image quality metrics compared to the baselines. For a fair comparison with ZeroNVS [43] and LucidDreamer [10], which are limited to square image inputs, we crop the generated novel views before computing the quantitative metrics. Figure 6 shows that scenes reconstructed using our method have less noise and distortion in occluded areas.
4.3 Ablation Study
Ablation on Video Diffusion Model. We evaluate the effectiveness of the design choices. As shown in Table 6, ray embedding achieves significant improvements in camera control accuracy and rendering quality compared to other camera encoding methods by introducing denser position encoding and optimizing geometric correspondences. Specifically, cross-frame consistency shows notable enhancement. Furthermore, when incorporating geometric constraints through Epipolar Attention, the system achieves optimal performance in terms of camera controllability and 3D consistency.
[…]
Raw Value Emb. 1.592 2.376 13.5% 81.854
Quaternion Emb. 1.457 2.394 13.1% 80.368
W/o Ray Emb. 2.683 3.145 15.7% 112.843
W/o Epipolar 1.125 2.235 12.2% 79.547
W/o Scale Align. 1.712 2.498 14.3% 81.679
Full model 1.014 2.218 4.2% 78.132
Table 6: Ablation study on the video diffusion model variants.
Figure 7: Ablation on the scale alignment. The video reconstruction results obtained based on scale-alignment training exhibit reduced artifacts while improving the 3D consistency of the scene.
Scale Alignment. As illustrated in Table 6 and Figure 7, the incorporation of dataset metric scale alignment significantly improves the pose control accuracy and 3D consistency of the generated video.
Video Latent-based Reconstruction. As demonstrated in Figure 8, incorporating generative priors and depth features in our method significantly enhances the fidelity of fine-grained geometric details in scene reconstruction. By introducing priors from video generative models, we reduced the difficulty of sparse-view reconstruction and achieved reasonable completion of unseen regions.
Figure 8: Ablation on the feed-forward Reconstruction Module.
Figure 9: Qualitative comparison of video results. Our method outperforms other video generation models by reducing the shot transition rate and ensuring smooth transitions, particularly when faced with large differences in input viewing angles.
4.4 Discussions
Existing commercial video generation models like Kling and GEN-3 Turbo face a critical challenge: significant differences in input viewing angles often result in unnatural camera transitions. As shown in Figure 9, our approach effectively addresses this issue by fine-tuning the video generation model on extensive 3D scene datasets and integrating specific 3D information during the inference process. This enables our model to produce videos with more continuous and natural camera transitions.
5 Conclusion
We present SpatialCrafter, a framework for scene reconstruction and novel view synthesis from sparse or single-view inputs. By leveraging video diffusion models to generate plausible additional observations, we effectively reduce the complexity of sparse-view scene reconstruction. Our key contributions include: (1) precise camera control via ray embeddings or depth-warped images with a trainable condition encoder; (2) a unified scale estimator solving the scale ambiguity in multi-dataset training; and (3) a hybrid Mamba-Transformer architecture that combines monocular depth priors with semantic features from video latent space to directly regress 3D Gaussian primitives. Experiments show our method outperforms existing methods, especially in single-view extrapolation and scenarios with little overlaps. Future work will focus on handling extreme complex geometries, and extending to dynamic scene reconstruction.
Appendix A Implementation Details
A.1 Training Details
Our video generation model is built upon SVD [3], an image-video diffusion model based on UNet. The training process incorporates a continuous-time noise scheduler (Karras et al., 2022), enabling the model to learn to gradually denoise high-variance Gaussian noise. Specifically, we define the noisy data as $x_{t}=x_{0}+n(t)$, where the added Gaussian noise $n(t)\sim\mathcal{N}(0,\sigma^{2}(t)I)$. The iterative denoising process corresponds to the probability flow ODE:
$dx=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(x;\sigma(t))dt,$
where $\nabla_{x}\log p(x;\sigma(t))$ is the score function parameterized by a neural network $D_{\theta}$. The training objective focuses on denoising score matching: $E\left[\|D_{\theta}(x_{0}+n;\sigma,c)-x_{0}\|^{2}_{2}\right]$, with $c$ representing the conditioning information. We follow the EDM framework, parameterizing the learnable denoiser as $D_{\theta}(x;\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(c_{\text{in}}(\sigma)x;c_{\text{noise}}(\sigma))$.
In our work, we use a relative camera system to convert all camera poses relative to the first frame located at the world origin. First, we train the model at a resolution of $320\times 512$ with a frame length of 25, for 50,000 iterations, and we perform center cropping on the input images to standardize the resolution. Subsequently, we adjust the model to a higher resolution of $576\times 1024$ and continue training for another 10,000 iterations, setting the learning rate to $51e-$ 5 and using the Adam optimizer with a 1,000-step warm-up. The training is conducted on 8 NVIDIA A800 GPUs with a batch size of 16. It is important to note that we only train the camera encoder, epipolar attention modules, and temporal attention while keeping the weights of other parts frozen. We first train on the Re10K and ACID datasets, with the sampling interval of video frames linearly increasing from 2 to 10. After that, we incorporate the DL3DV dataset, randomly sampling the video frame intervals between $[2,10]$ to enable the model to adapt to different motion speeds. We choose Lightning as the training framework, utilizing mixed precision fp16 and DeepSpeed ZeRO-2 to enhance efficiency. During inference, we implement the DDIM sampler [49] alongside classifier-free guidance [21] to enhance performance.
A.2 Evaluation Details
Video Generation. When using COLMAP to extract the camera pose of the generated video, the reliability of the obtained pose is often low. Therefore, we employ DUSt3R for a more robust pose estimation. To compare with previous methods, we transform the camera coordinates of the estimated poses to be relative to the first frame and normalize the scale using the furthest frame. We then calculate the rotation distance relative to the ground truth rotation matrices of each generated novel view sequence, expressed as $R_{\text{dist}}=\sum_{i=1}^{n}\arccos\left(\frac{\text{tr}(R_{\text{gen}}^{i}R_{\text{gt}}^{i^{T}})-1}{2}\right)$. We also compute the $T_{dist}=\sum_{i=1}^{n}\|T_{gt}^{i}-T_{gen}^{i}\|_{2}$, which measures the Euclidean distance between ground truth and estimated translations across $n$ video frames.
Sparse View Reconstruction. To ensure a fair comparison with feed-forward baselines [8, 9], we conducted experiments on $256\times 256$ resolutions. To comprehensively evaluate our method, we compared the results of input views under small and large overlap ratios. We employed RoMa to determine the overlap ratio of image pairs. First, we obtained dense matches from the first image to the second and vice versa, considering pixels with matching scores exceeding 0.005 as valid. Next, we calculated the overlap ratio for each image pair by dividing the number of valid matched pixels by the total number of pixels, with the final overlap ratio defined as the minimum of the two directional calculations. Additionally, to compare with optimization-based reconstruction baselines [14, 29], we used only 2 ground truth training images per scene and evaluated using 12 views. The evaluation metrics included PSNR, SSIM, and LPIPS.
Appendix B Discussions
B.1 Comparison with Very Recent Methods
ViewCrafter. ViewCrafter [71] leverages point cloud rendering as a condition for fine-tuning video diffusion models. However, constrained by Dust3R’s capabilities, the approach may encounter challenges in synthesizing novel views with extensive view range transformations, particularly when generating front-view images from limited back-view inputs. Moreover, in scenarios where DUSt3R fails—such as incorrectly predicting planar geometries—ViewCrafter’s performance can be significantly compromised. In contrast, our proposed generation module decouples from the reconstruction module, effectively mitigating these limitations and enhancing view synthesis robustness.
DimensionX. DimensionX [51] achieves promising results by fine-tuning respective LORA models for different motion patterns. However, this method has some limitations: first, it cannot generate more complex motion patterns; second, it lacks effective control over the speed and amplitude of motion. In contrast, our method introduces camera embedding as the conditional signal, which effectively solves the above problems, thereby enabling the generation of more complex motion patterns and providing precise control over the speed and amplitude of motion.
ReconX. ReconX [33] initially constructs a global point cloud representation and encodes it into a contextual space, serving as 3D geometric conditional information for video diffusion models. However, the method’s performance critically depends on the geometric reconstruction accuracy of the underlying DUST3R algorithm. When DUST3R fails to precisely align and reconstruct geometric structures, the point cloud information injected into the video diffusion model may introduce significant geometric distortions, leading to suboptimal generation results. More importantly, the current approach is limited to multi-view interpolation scenarios, lacking an effective solution for single-view extrapolation tasks
CAT3D. CAT3D [16] extends the text-to-image (T2I) generation model by a large-scale data-driven method to learn a 3D prior of the scene, which usually requires a large amount of training time. In addition, the generation stage requires rendering a large number of frames (about 100 frames) to reconstruct the scene. We recognize that video generation models trained on large video datasets embed inherent scene prior information, and therefore propose a new method based on fine-tuning with a video diffusion model. Our method speeds up convergence and retains the generative prior of the pre-trained video model, thereby enhancing the model’s generalization ability. By introducing explicit geometric constraints, we further improve the 3D consistency of the generated scenes. Notably, during the inference stage, we utilize the geometric prior for initialization and reconstruction, greatly reducing the number of images required and the optimization time.
Appendix C More Qualitative Results
C.1 More Generated Video Results
As shown in Figure 14, we present a series of generated video sequences covering different scenarios, including indoor environments, outdoor landscapes, and object-centered scenes. Our method shows excellent stability in generating video frame sequences with high 3D consistency, and can effectively predict the plausible appearance details of previously unseen regions. In addition, Figure 12 demonstrates the view interpolation ability of our video generation model using sparse input views, and the performance remains stable even when the input view angles differ greatly.
C.2 Basic Camera Trajectories
Our method demonstrates promising generalization ability in camera-controllable video generation, especially for common motion patterns such as zoom-in, zoom-out, and panning. As shown in Figure 10, our method can generate stable videos with basic camera trajectories using a large-scale training dataset that contains a diverse and complex distribution of camera motion patterns. In addition, in Figure 11, we also show the control ability of our method for different motion amplitudes.
C.3 Single View Exploration Results
Figure 13 illustrates that scene reconstruction from sparse input views often suffers from severe occlusions and geometric gaps due to limited observation angles. By implementing a multi-view observation strategy using video diffusion models, we can progressively fill in these missing regions, thereby enhancing the reconstruction’s completeness and geometric detail. As we continuously increase and synthesize observation perspectives, the scene’s geometric structure becomes more complete, significantly improving the quality and accuracy of the 3D reconstruction. In Figure 11, we show that by controlling the camera parameters, we can directly control the magnitude of the movement, which is general in different scenarios.
C.4 More 3D Reconstruction Results
Figure 16 presents additional reconstruction and rendering results based on generated video frames, demonstrating that our method effectively achieves high-quality reconstructions in indoor, outdoor, and object-centered scenes.
Appendix D Limitations
Our method encounters several challenges in video generation and scene reconstruction. First, existing video diffusion models suffer from significant inference efficiency bottlenecks, particularly when processing large-scale scene reconstruction tasks requiring numerous video frames, leading to prohibitively high computational costs. Second, current video diffusion models are constrained in the number of generated video frames, which impedes comprehensive scene reconstruction for complex scenarios. Lastly, the lack of effective user interaction mechanisms represents a substantial limitation, as users cannot exert fine-grained control over the generation process and results, restricting the model’s adaptability in practical applications.
Figure 10: Visualization of our method for generating videos with basic motion trajectories.
Figure 11: Visualization of our method for controlling different motion amplitudes.
Figure 12: More Visualization of interpolated video based on sparse view input.
Figure 13: Visualization of exploration and reconstruction based on single-view input.
Figure 14: More video generation results with high 3D consistency based on single-view input.
Figure 15: Consistent frames generated by our approach. Our method demonstrates strong generalization capability across diverse scenarios, including the DL3DV [31], Tanks-and-Temples benchmark [27], and synthetic images from generative models, while maintaining geometric coherence and structural fidelity.
Figure 16: More video generation and reconstruction results with high 3D consistency based on single-view input.

### v2

@@
 Existing techniques, though progressed, rely on dense multi-view observations, restricting their application.
-This work takes on the challenge of reconstructing photorealistic 3D scenes from sparse or single-view inputs.
+We tackle the task of reconstructing photorealistic 3D scenes from only one or a few input views.
 We introduce SpatialCrafter, a framework that leverages the rich knowledge in video diffusion models to generate plausible additional observations, thereby alleviating reconstruction ambiguity.
@@
 Extensive experiments show our method enhances sparse view reconstruction and restores the realistic appearance of 3D scenes.
+Project page: https://franklinz233.github.io/projects/spatialcrafter/.
 Figure 1: Overview.
@@
 More visualization results can be found in the supplementary material.
+footnotetext: * Corresponding author
 1 Introduction
@@
 Therefore, this paper focuses on a more practically valuable challenge: how to achieve high-quality 3D scene reconstruction and synthesize realistic novel views from sparse or even single-view observation.
-The complexity of this problem stems from three main aspects: first, the diversity and complexity of real-world scene make precise modeling extremely challenging; second, occluded regions in sparse viewpoints are difficult to recover; and finally, limited geometric cues constrain the accuracy of reconstruction.
-Compared to traditional methods that depend on numerous viewpoints, achieving high-quality reconstruction from sparse views requires stronger prior knowledge and more efficient reasoning capabilities, which is the core challenge of this problem.
+This problem is hard because (1) real-world scenes are diverse and complex, (2) occluded regions are unseen in sparse captures, and (3) geometric cues are too few to constrain reconstruction.
 To address these challenges, some methods [29, 88, 26, 40, 17, 84] constrain the optimization process by introducing regularization terms, but these methods need to be optimized for each scene, are not generalizable, and are difficult to cope with complex scenarios.
 Recently, some feed-forward methods [7, 8, 60, 50, 22, 55, 79, 80] have achieved generalizable sparse view scene reconstruction by directly predicting 3D Gaussian primitives for each pixel.
-However, such methods produce severe artifacts and implausible images when reconstructing occluded regions and applying them to the extrapolation setting.
+However, such methods produce severe artifacts when reconstructing occluded regions and applying them to the extrapolation setting.
 To address this challenge, some methods [35, 64, 45, 65] have introduced the priors from the diffusion model to achieve sparse or single-view NVS, but they perform poorly on scene-level data and lack precise pose control and 3D consistency.
@@
 Specifically, we first focus on enhancing precise camera control and 3D consistency in generated videos.
-To achieve this, we parameterize camera settings using ray embeddings or metric depth-warped images and incorporate them into the video diffusion model via a trainable camera encoder.
+To achieve this, we parameterize camera settings using ray embeddings [78] or metric depth-warped images [77] and incorporate them into the video diffusion model via a trainable camera encoder.
 Furthermore, we propose an epipolar attention mechanism that improves 3D consistency between video frames through explicit geometric constraints.
@@
 Figure 2: Overview of our pipeline.
-Our generative reconstruction pipeline consists of two parts: camera-controlled video generation, and then reconstruction.
+Our generative reconstruction pipeline consists of two parts: camera-controlled video generation, and then video-based reconstruction.
 First, we set the exploration path based on the input views.
-In the video generation module, ray embedding is used to parameterize the camera information, and a cross-attention mechanism with epipolar constraints is introduced to improve the 3D consistency of the generated video.
-The robust 3D scene reconstruction pipeline integrates monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives through a feed-forward manner.
+In the video generation module, ray embeddings are used to parameterize the camera trajectories, and a cross-attention mechanism with epipolar constraints is introduced to improve the 3D consistency of the generated video.
+The video-based scene reconstruction pipeline integrates monocular depth priors with semantic features extracted from the video latent space, and directly regress 3D Gaussian primitives via a feed-forward manner.
 2.1 Sparse-View Scene Reconstruction
@@
 To address this problem, some approaches [29, 17, 88, 71] focus on constraining the optimization process of 3D representations via manually designed regularization.
-Some other methods [23, 8, 9, 52, 65, 68] are trained on large-scale datasets and directly predict the 3D representation in a feed-forward manner.
-Recently, some approaches [14, 33, 48, 30] based on end-to-end 3D reconstruction models [56, 28], have achieved efficient and high-quality 3D reconstruction.
-However, these methods cannot handle occluded areas, and when reconstruction fails (e.g., in the case of large viewpoint changes), the results of subsequent novel view synthesis can be significantly compromised.
+Some other methods [23, 7, 8, 54, 69, 72, 58, 83] are trained on large-scale datasets and directly predict the 3D representation in a feed-forward manner.
+Recently, some approaches [13, 50] based on end-to-end 3D reconstruction models [60, 28, 7], have achieved efficient and high-quality 3D reconstruction.
+However, these methods cannot handle occluded regions, and reconstruction failures (e.g., under large viewpoint changes) can severely affect the quality of novel view synthesis.
+More relevant to our work, some works [57, 34, 31] have employed priors from video diffusion models to enhance the performance of sparse-view reconstruction.
 2.2 Diffusion-based Novel View Synthesis
@@
 Some approaches [18, 53] employs motion-specific training for predefined camera movements, though this framework struggles with complex trajectory synthesis.
-Some methods [60, 20, 63] enable more complex camera control by parameterizing the camera trajectory and injecting it into a pre-trained video diffusion model via a trainable camera encoder.
-MotionCtrl [60] controls camera and object motions independently through learnable adapters, while CameraCtrl [20] introduces Plücker embeddings for accurate 6-DoF camera pose representation.
-AC3D [1] optimizes the pose adjustment scheme to accelerate convergence and limits the injection of camera features to specific locations.
-ViewCrafter [71] and ReconX [33] pushes this direction by enabling pixel-accurate view compositing through point-based rendering formulation.
+Some methods [64, 19, 20, 67, 1] enable more complex camera control by parameterizing the camera trajectory and injecting it into a pre-trained video diffusion model via a trainable camera encoder.
+Another line of work [77, 34, 74, 43] advances this direction by enabling pixel-accurate view compositing through a point-based rendering approach.
 3 Method
@@
 Next, we introduce ourscene reconstruction pipeline (Section 3.2), which aims to perform stable 3D reconstruction based on the generated video latents.
-These innovations collectively improve the accuracy and consistency of 3D reconstructions from generated video.
 3.1 Camera-Conditioned Video Generation
 Scale Alignment.
-While recent methods [60, 20] have achieved promising results in novel view synthesis, their reliance on ground-truth camera extrinsic from Re10K [79] during inference inherently limits their applicability to in-the-wild scenarios.
-Moreover, the metric-scaled poses in Re10K create alignment challenges when jointly training with other datasets (e.g., DL3DV [31]) where poses are estimated using COLMAP, primarily due to scale ambiguity issues.
-To address this issue, we introduced a metric depth estimation model [5] to achieve scale alignment of camera poses between different datasets.
-During the data preprocessing stage, we align the metric 3D point cloud predicted by the model with the sparse point cloud generated using COLMAP to calculate the scale factor, thereby obtaining absolute-scale camera extrinsics.
+To address this challenge, our approach first employs VGGT [58] to estimate initial camera parameters, represented by the extrinsic matrix $\mathbf{P}=[\mathbf{R}|\mathbf{T}]$, and a corresponding depth map $d_{v}$ for each video frame.
+To establish a consistent scale across these datasets, we leverage the Metric3D [73] to infer a canonical metric depth map, denoted $d_{m}$.
+We then align our initial depth prediction, $d_{v}$, to the canonical scale by calculating a scale factor $s$, defined as the ratio of the inter-percentile ranges (IPR) of the two depth maps: $s=\operatorname{IPR}_{0.8,0.2}(d_{v})/\operatorname{IPR}_{0.8,0.2}(d_{m})$.
+This scale factor is then used to adjust the translation component of the camera extrinsics, resulting in $\mathbf{P}_{scaled}=[\mathbf{R}|s\cdot\mathbf{T}]$.
+This procedure yields camera extrinsics with a consistent absolute scale, effectively resolving alignment issues when fusing data from different sources.
 Camera Injection.
 Inspired by recent methods [19, 78, 49], we chose to use ray embeddings [19] or the depth warping frames [77] to represent the camera information.
-This over-parameterized approach [72] provides more accurate pose information to the video generator than other methods.
 Specifically, for a ray defined by the origin $o\in\mathbb{R}^{3}$ and the normalized direction $d\in\mathbb{R}^{3}$, we can represent it as $(o\times d,d)\in\mathbb{R}^{6}$.
@@
 The training objective is:
-$\mathcal{L}=\mathbb{E}_{z,z_{0},\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},t,\phi({c}))\right\|_{2}^{2}\right]$ (1)
-where $\boldsymbol{P}$ represents the ray embeddings, $z_{0}$ denotes the latent embeddings of the reference frame, and $t$ indicates the timestamps.
+$\mathcal{L}=\mathbb{E}_{z,z_{0},\epsilon,{C},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},t,\phi({C}))\right\|_{2}^{2}\right]$ (1)
+where $\phi({C})$ represents the camera conditional encodings, $z_{0}$ denotes the latent embeddings of the reference frame, and $t$ indicates the timestamps.
 Epipolar Feature Aggregation.
@@
 The training objective is:
-$\mathcal{L}=\mathbb{E}_{z,z_{0},z_{n}\epsilon,\boldsymbol{P},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},z_{n},t,\boldsymbol{P})\right\|_{2}^{2}\right]$ (2)
+$\mathcal{L}=\mathbb{E}_{z,z_{0},z_{n}\epsilon,{C},t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t};z_{0},z_{n},t,\phi({C})))\right\|_{2}^{2}\right]$ (2)
 where $z_{n}$ and $z_{0}$ are the latent embeddings of the final and initial frames, respectively.
-During the inference phase, we use DUSt3R to estimate the extrinsic and intrinsic parameters for the input views.
+During the inference phase, we use VGGT [58] to estimate the extrinsic and intrinsic parameters for the input views.
 Figure 3: Qualitative comparison of generated videos.
@@
 To address this, a powerful reconstruction module specifically designed to handle the generated videos is required.
+Latent Feature Fusion.
+Given input video latents $\mathbf{z}\in\mathbb{R}^{T\times H\times W\times C}$ and camera poses encoded as ray embeddings $\mathbf{p}\in\mathbb{R}^{T\times H\times W\times 6}$, we transform them into token sequences through a patchification process.
+Specifically, we apply spatial patchification to latent features to obtain latent tokens $\mathbf{z}_{t}\in\mathbb{R}^{N\times d_{z}}$, while ray embeddings undergo 3D-patchification to produce pose tokens $\mathbf{p}_{t}\in\mathbb{R}^{N\times d_{p}}$.
+To incorporate explicit geometric guidance, we leverage monocular depth estimation within our pipeline.
+For each input RGB video frame, we estimate dense depth maps $\mathbf{D}\in\mathbb{R}^{T\times H\times W\times 1}$, which are processed by a dedicated depth encoder to yield depth tokens $\mathbf{d}_{t}\in\mathbb{R}^{N\times d_{d}}$.
+The three token sets are channel-wise concatenated to form a unified representation $\mathbf{x}=[\mathbf{z}_{t};\mathbf{p}_{t};\mathbf{d}_{t}]\in\mathbb{R}^{N\times(d_{z}+d_{p}+d_{d})}$, which is then linearly projected to a lower-dimensional space $\mathbf{x}^{\prime}\in\mathbb{R}^{N\times d}$ before being fed into a sequence of mamba and transformer blocks.
+Mamba [16] achieves the same token sequence processing functionality as Transformer, but reduces computational complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L)$, making it particularly suitable for dense reconstruction tasks.
+In our Mamba blocks, we implement bi-directional scanning across the token sequence.
+First, we compute state parameters $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times d}$ from the input tokens using a linear projection.
+Then, we execute the State Space Model in both forward ($\mathbf{y}_{f}$) and backward ($\mathbf{y}_{b}$) directions.
+The final output of each Mamba block is computed as $\mathbf{y}=\mathbf{y}_{f}+\mathbf{y}_{b}$ followed by a final linear transformation.
+Gaussian Decoding.
+We design a lightweight decoder module that efficiently transforms output feature tokens into per-pixel Gaussian parameters.
+The decoding module consists of 3D-DeConv layers, which generates refined Gaussian feature map $G\in\mathbb{R}^{(T\times H\times W)\times 12}$.
+This 12-channel representation precisely encodes the complete set of Gaussian parameters: RGB color (3 channels), scale factors (3 channels), rotation quaternion (4 channels), opacity (1 channel), and ray distance (1 channel).
+The final output of the model is the merge of 3D Gaussians from all input video frames.
+Training Objective. During training, we render images from predicted Gaussians using randomly selected supervision views. Our approach employs a composite loss that integrates three components:
+$\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mse}+\lambda_{2}\mathcal{L}_{perc}+\lambda_{3}\mathcal{L}_{depth},$ (3)
+where $\mathcal{L}_{mse}$ represents pixel-wise mean squared error, $\mathcal{L}_{perc}$ denotes perceptual loss, and $\mathcal{L}_{depth}$ enforces depth consistency.
+This formulation jointly optimizes for photometric accuracy, high-level perceptual fidelity, and geometric coherence.
 […]
@@
 Ours 18.21 171.52 0.063 0.134 0.225 17.02 0.537
-Table 1: Quantitative comparison to the camera conditioned video generation method on RealEstate10K, DL3DV, and Tanks-and-Temples dataset.
-Latent Feature Fusion.
-Given input video latents $\mathbf{z}\in\mathbb{R}^{T\times H\times W\times C}$ and camera poses encoded as ray embeddings $\mathbf{p}\in\mathbb{R}^{T\times H\times W\times 6}$, we transform them into token sequences through a patchification process.
-Specifically, we apply spatial patchification to latent features to obtain latent tokens $\mathbf{z}_{t}\in\mathbb{R}^{N\times d_{z}}$, while ray embeddings undergo 3D-patchification to produce pose tokens $\mathbf{p}_{t}\in\mathbb{R}^{N\times d_{p}}$.
-To incorporate explicit geometric guidance, we leverage monocular depth estimation within our pipeline.
-For each input RGB video frame, we estimate dense depth maps $\mathbf{D}\in\mathbb{R}^{T\times H\times W\times 1}$, which are processed by a dedicated depth encoder to yield depth tokens $\mathbf{d}_{t}\in\mathbb{R}^{N\times d_{d}}$.
-The three token sets are channel-wise concatenated to form a unified representation $\mathbf{x}=[\mathbf{z}_{t};\mathbf{p}_{t};\mathbf{d}_{t}]\in\mathbb{R}^{N\times(d_{z}+d_{p}+d_{d})}$, which is then linearly projected to a lower-dimensional space $\mathbf{x}^{\prime}\in\mathbb{R}^{N\times d}$ before being fed into a sequence of mamba and transformer blocks.
-Mamba [17] achieves the same token sequence processing functionality as Transformer, but reduces computational complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L)$, making it particularly suitable for dense reconstruction tasks.
-In our Mamba blocks, we implement bi-directional scanning across the token sequence.
-First, we compute state parameters $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{N\times d}$ from the input tokens using a linear projection.
-Then, we execute the State Space Model (SSM) in both forward ($\mathbf{y}_{f}$) and backward ($\mathbf{y}_{b}$) directions.
-The final output of each Mamba block is computed as $\mathbf{y}=\mathbf{y}_{f}+\mathbf{y}_{b}$ followed by a final linear transformation.
-Gaussian Decoding.
-We design a lightweight decoder module that efficiently transforms output feature tokens into per-pixel Gaussian parameters.
-The decoding module consists of 3D-DeConv layers, which generates refined Gaussian feature map $G\in\mathbb{R}^{(T\times H\times W)\times 12}$.
-This 12-channel representation precisely encodes the complete set of Gaussian parameters: RGB color (3 channels), scale factors (3 channels), rotation quaternion (4 channels), opacity (1 channel), and ray distance (1 channel).
-The final output of the model is the merge of 3D Gaussians from all input video frames.
-Training Objective. During training, we render images from predicted Gaussians using randomly selected supervision views. Our approach employs a composite loss that integrates three components:
-$\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mse}+\lambda_{2}\mathcal{L}_{perc}+\lambda_{3}\mathcal{L}_{depth},$ (3)
-where $\mathcal{L}_{mse}$ represents pixel-wise mean squared error, $\mathcal{L}_{perc}$ denotes perceptual loss, and $\mathcal{L}_{depth}$ enforces depth consistency.
-This formulation jointly optimizes for photometric accuracy, high-level perceptual fidelity, and geometric coherence.
+Table 1: Quantitative comparison to the camera conditioned video generation method on RealEstate10K [86], DL3DV [32], and Tanks-and-Temples [27] dataset.
 4 Experiment
@@
 Table 2: Quantitative comparison of sparse view 3D reconstruction on Out-of-Domain Datasets.
+4.2 Comparisons
+4.2.1 Controllable Video Generation
+Benchmark and Metrics.
+We evaluate our video generation model against three baselines [64, 19, 77] on three benchmark datasets: RealEstate10K [86] (500 randomly selected videos with first frame as image condition and subsequent frames at stride 3 for pose guidance), DL3DV-140 [32] (500 video clips at stride 2), and Tanks-and-Temples [27] (100 sequences at stride 4 from 14 scenes with COLMAP-annotated poses) for out-of-domain generalization testing.
+Metrics.
+Our evaluation framework employs multiple metrics to assess performance: visual quality measured through Fréchet Video Distance (FVD) and Fréchet Inception Distance (FID); camera control precision quantified by rotation error ($R_{err}$) and translation error ($T_{err}$) computed from camera poses extracted via COLMAP and normalized relative to the first frame; and visual consistency evaluated using PSNR, SSIM, and LPIPS [81] between generated and ground-truth views.
+To ensure fair comparison across methods with varying output capabilities, we restrict visual similarity assessment to the first 14 frames, as generated content typically diverges progressively from the conditional single-view input as the scene extends.
 RealEstate10K ACID
@@
 Table 3: Quantitative comparison of two-view sparse view reconstruction methods on RealEstate10K [86] and ACID [33] dataset.
-4.2 Comparisons
-4.2.1 Controllable Video Generation
-Benchmark and Metrics.
-We evaluate our video generation model against three baselines [60, 20, 71] on three benchmark datasets: RealEstate10K [79] (500 randomly selected videos with first frame as image condition and subsequent frames at stride 3 for pose guidance), DL3DV-140 [31] (500 video clips at stride 2), and Tanks-and-Temples [27] (100 sequences at stride 4 from 14 scenes with COLMAP-annotated poses) for out-of-domain generalization testing.
-Metrics.
-Our evaluation framework employs multiple metrics to assess performance: visual quality measured through Fréchet Video Distance (FVD) and Fréchet Inception Distance (FID); camera control precision quantified by rotation error ($R_{err}$) and translation error ($T_{err}$) computed from camera poses extracted via COLMAP and normalized relative to the first frame; and visual consistency evaluated using PSNR, SSIM, and LPIPS [75] between generated and ground-truth views.
-To ensure fair comparison across methods with varying output capabilities, we restrict visual similarity assessment to the first 14 frames, as generated content typically diverges progressively from the conditional single-view input as the scene extends.
 Comparison.
@@
 Ours 22.514 0.784 0.213 27.411 0.913 0.109
-Table 4: Quantitative comparison on the RealEstate10K dataset.
+Table 4: Quantitative comparison on the RealEstate10K [86] dataset.
 Small and Large refer to input images with low and high overlap ratios, respectively.
 Greater overlap means greater temporal coherence between adjacent images.
+Tanks-and-Temples DL3DV
+[…]
+LucidDreamer [9] 14.552 0.364 0.415 15.126 0.452 0.431
+ZeroNVS [45] 14.734 0.381 0.480 15.163 0.464 0.466
+MotionCtrl [64] 15.322 0.426 0.404 16.889 0.528 0.392
+ViewCrafter [77] 21.286 0.654 0.193 21.373 0.686 0.244
+Ours 22.331 0.742 0.213 22.782 0.791 0.202
+Table 5: Quantitative comparison of single-view novel view synthesis on DL3DV [32] and Tank-and-Temples [27] benchmarks.
 4.2.2 Sparse View Reconstruction
@@
 Figure 6: Qualitative comparisons of novel views rendered from scenes reconstructed using other 3D generation methods.
-Tanks-and-Temples DL3DV
-[…]
-LucidDreamer [10] 14.552 0.364 0.415 15.126 0.452 0.431
-ZeroNVS [43] 14.734 0.381 0.480 15.163 0.464 0.466
-MotionCtrl [60] 15.322 0.426 0.404 16.889 0.528 0.392
-ViewCrafter [71] 21.286 0.654 0.193 21.373 0.686 0.244
-Ours 22.331 0.742 0.213 22.782 0.791 0.202
-Table 5: Quantitative comparison of single-view novel view synthesis on DL3DV [31] and Tank-and-Temples [27] benchmarks.
+Figure 7: Additional visualizations of generated videos and their corresponding 3DGS representations.
 4.2.3 Single View 3D Generation
@@
 Figure 6 shows that scenes reconstructed using our method have less noise and distortion in occluded areas.
+For more visualizations, see Figure 7.
 4.3 Ablation Study
@@
 Figure 9: Ablation on the feed-forward Reconstruction Module.
-Figure 9: Qualitative comparison of video results.
-Our method outperforms other video generation models by reducing the shot transition rate and ensuring smooth transitions, particularly when faced with large differences in input viewing angles.
-4.4 Discussions
-Existing commercial video generation models like Kling and GEN-3 Turbo face a critical challenge: significant differences in input viewing angles often result in unnatural camera transitions.
-As shown in Figure 9, our approach effectively addresses this issue by fine-tuning the video generation model on extensive 3D scene datasets and integrating specific 3D information during the inference process.
-This enables our model to produce videos with more continuous and natural camera transitions.
 5 Conclusion
@@
 Experiments show our method outperforms existing methods, especially in single-view extrapolation and scenarios with little overlaps.
-Future work will focus on handling extreme complex geometries, and extending to dynamic scene reconstruction.
-Appendix A Implementation Details
-A.1 Training Details
-Our video generation model is built upon SVD [3], an image-video diffusion model based on UNet.
-The training process incorporates a continuous-time noise scheduler (Karras et al., 2022), enabling the model to learn to gradually denoise high-variance Gaussian noise.
-Specifically, we define the noisy data as $x_{t}=x_{0}+n(t)$, where the added Gaussian noise $n(t)\sim\mathcal{N}(0,\sigma^{2}(t)I)$.
-The iterative denoising process corresponds to the probability flow ODE:
-$dx=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(x;\sigma(t))dt,$
-where $\nabla_{x}\log p(x;\sigma(t))$ is the score function parameterized by a neural network $D_{\theta}$.
-The training objective focuses on denoising score matching: $E\left[\|D_{\theta}(x_{0}+n;\sigma,c)-x_{0}\|^{2}_{2}\right]$, with $c$ representing the conditioning information.
-We follow the EDM framework, parameterizing the learnable denoiser as $D_{\theta}(x;\sigma)=c_{\text{skip}}(\sigma)x+c_{\text{out}}(\sigma)F_{\theta}(c_{\text{in}}(\sigma)x;c_{\text{noise}}(\sigma))$.
-In our work, we use a relative camera system to convert all camera poses relative to the first frame located at the world origin.
-First, we train the model at a resolution of $320\times 512$ with a frame length of 25, for 50,000 iterations, and we perform center cropping on the input images to standardize the resolution.
-Subsequently, we adjust the model to a higher resolution of $576\times 1024$ and continue training for another 10,000 iterations, setting the learning rate to $51e-$ 5 and using the Adam optimizer with a 1,000-step warm-up.
-The training is conducted on 8 NVIDIA A800 GPUs with a batch size of 16.
-It is important to note that we only train the camera encoder, epipolar attention modules, and temporal attention while keeping the weights of other parts frozen.
-We first train on the Re10K and ACID datasets, with the sampling interval of video frames linearly increasing from 2 to 10.
-After that, we incorporate the DL3DV dataset, randomly sampling the video frame intervals between $[2,10]$ to enable the model to adapt to different motion speeds.
-We choose Lightning as the training framework, utilizing mixed precision fp16 and DeepSpeed ZeRO-2 to enhance efficiency.
-During inference, we implement the DDIM sampler [49] alongside classifier-free guidance [21] to enhance performance.
-A.2 Evaluation Details
-Video Generation.
-When using COLMAP to extract the camera pose of the generated video, the reliability of the obtained pose is often low.
-Therefore, we employ DUSt3R for a more robust pose estimation.
-To compare with previous methods, we transform the camera coordinates of the estimated poses to be relative to the first frame and normalize the scale using the furthest frame.
-We then calculate the rotation distance relative to the ground truth rotation matrices of each generated novel view sequence, expressed as $R_{\text{dist}}=\sum_{i=1}^{n}\arccos\left(\frac{\text{tr}(R_{\text{gen}}^{i}R_{\text{gt}}^{i^{T}})-1}{2}\right)$.
-We also compute the $T_{dist}=\sum_{i=1}^{n}\|T_{gt}^{i}-T_{gen}^{i}\|_{2}$, which measures the Euclidean distance between ground truth and estimated translations across $n$ video frames.
-Sparse View Reconstruction.
-To ensure a fair comparison with feed-forward baselines [8, 9], we conducted experiments on $256\times 256$ resolutions.
-To comprehensively evaluate our method, we compared the results of input views under small and large overlap ratios.
-We employed RoMa to determine the overlap ratio of image pairs.
-First, we obtained dense matches from the first image to the second and vice versa, considering pixels with matching scores exceeding 0.005 as valid.
-Next, we calculated the overlap ratio for each image pair by dividing the number of valid matched pixels by the total number of pixels, with the final overlap ratio defined as the minimum of the two directional calculations.
-Additionally, to compare with optimization-based reconstruction baselines [14, 29], we used only 2 ground truth training images per scene and evaluated using 12 views.
-The evaluation metrics included PSNR, SSIM, and LPIPS.
-Appendix B Discussions
-B.1 Comparison with Very Recent Methods
-ViewCrafter.
-ViewCrafter [71] leverages point cloud rendering as a condition for fine-tuning video diffusion models.
-However, constrained by Dust3R’s capabilities, the approach may encounter challenges in synthesizing novel views with extensive view range transformations, particularly when generating front-view images from limited back-view inputs.
-Moreover, in scenarios where DUSt3R fails—such as incorrectly predicting planar geometries—ViewCrafter’s performance can be significantly compromised.
-In contrast, our proposed generation module decouples from the reconstruction module, effectively mitigating these limitations and enhancing view synthesis robustness.
-DimensionX.
-DimensionX [51] achieves promising results by fine-tuning respective LORA models for different motion patterns.
-However, this method has some limitations: first, it cannot generate more complex motion patterns; second, it lacks effective control over the speed and amplitude of motion.
-In contrast, our method introduces camera embedding as the conditional signal, which effectively solves the above problems, thereby enabling the generation of more complex motion patterns and providing precise control over the speed and amplitude of motion.
-ReconX.
-ReconX [33] initially constructs a global point cloud representation and encodes it into a contextual space, serving as 3D geometric conditional information for video diffusion models.
-However, the method’s performance critically depends on the geometric reconstruction accuracy of the underlying DUST3R algorithm.
-When DUST3R fails to precisely align and reconstruct geometric structures, the point cloud information injected into the video diffusion model may introduce significant geometric distortions, leading to suboptimal generation results.
-More importantly, the current approach is limited to multi-view interpolation scenarios, lacking an effective solution for single-view extrapolation tasks
-CAT3D.
-CAT3D [16] extends the text-to-image (T2I) generation model by a large-scale data-driven method to learn a 3D prior of the scene, which usually requires a large amount of training time.
-In addition, the generation stage requires rendering a large number of frames (about 100 frames) to reconstruct the scene.
-We recognize that video generation models trained on large video datasets embed inherent scene prior information, and therefore propose a new method based on fine-tuning with a video diffusion model.
-Our method speeds up convergence and retains the generative prior of the pre-trained video model, thereby enhancing the model’s generalization ability.
-By introducing explicit geometric constraints, we further improve the 3D consistency of the generated scenes.
-Notably, during the inference stage, we utilize the geometric prior for initialization and reconstruction, greatly reducing the number of images required and the optimization time.
-Appendix C More Qualitative Results
-C.1 More Generated Video Results
-As shown in Figure 14, we present a series of generated video sequences covering different scenarios, including indoor environments, outdoor landscapes, and object-centered scenes.
-Our method shows excellent stability in generating video frame sequences with high 3D consistency, and can effectively predict the plausible appearance details of previously unseen regions.
-In addition, Figure 12 demonstrates the view interpolation ability of our video generation model using sparse input views, and the performance remains stable even when the input view angles differ greatly.
-C.2 Basic Camera Trajectories
-Our method demonstrates promising generalization ability in camera-controllable video generation, especially for common motion patterns such as zoom-in, zoom-out, and panning.
-As shown in Figure 10, our method can generate stable videos with basic camera trajectories using a large-scale training dataset that contains a diverse and complex distribution of camera motion patterns.
-In addition, in Figure 11, we also show the control ability of our method for different motion amplitudes.
-C.3 Single View Exploration Results
-Figure 13 illustrates that scene reconstruction from sparse input views often suffers from severe occlusions and geometric gaps due to limited observation angles.
-By implementing a multi-view observation strategy using video diffusion models, we can progressively fill in these missing regions, thereby enhancing the reconstruction’s completeness and geometric detail.
-As we continuously increase and synthesize observation perspectives, the scene’s geometric structure becomes more complete, significantly improving the quality and accuracy of the 3D reconstruction.
-In Figure 11, we show that by controlling the camera parameters, we can directly control the magnitude of the movement, which is general in different scenarios.
-C.4 More 3D Reconstruction Results
-Figure 16 presents additional reconstruction and rendering results based on generated video frames, demonstrating that our method effectively achieves high-quality reconstructions in indoor, outdoor, and object-centered scenes.
-Appendix D Limitations
-Our method encounters several challenges in video generation and scene reconstruction.
-First, existing video diffusion models suffer from significant inference efficiency bottlenecks, particularly when processing large-scale scene reconstruction tasks requiring numerous video frames, leading to prohibitively high computational costs.
-Second, current video diffusion models are constrained in the number of generated video frames, which impedes comprehensive scene reconstruction for complex scenarios.
-Lastly, the lack of effective user interaction mechanisms represents a substantial limitation, as users cannot exert fine-grained control over the generation process and results, restricting the model’s adaptability in practical applications.
-Figure 10: Visualization of our method for generating videos with basic motion trajectories.
-Figure 11: Visualization of our method for controlling different motion amplitudes.
-Figure 12: More Visualization of interpolated video based on sparse view input.
-Figure 13: Visualization of exploration and reconstruction based on single-view input.
-Figure 14: More video generation results with high 3D consistency based on single-view input.
-Figure 15: Consistent frames generated by our approach.
-Our method demonstrates strong generalization capability across diverse scenarios, including the DL3DV [31], Tanks-and-Temples benchmark [27], and synthetic images from generative models, while maintaining geometric coherence and structural fidelity.
-Figure 16: More video generation and reconstruction results with high 3D consistency based on single-view input.
+Future work will focus on extending to dynamic scene.